A CD-ROM drive in a computer spins the 12-cm-diameter disks at 9900 rpm.

1) What is a disk's period (in s) ?

2) What is a disk's frequency (in rev/s)?

3) What is the acceleration in units of g that this speck of dust experiences?
Answer #1

Angular speed = 9900 rpm = 9900×2×pi/60 rad/s

W = 1036.725 rad/s

1)Time period T = 2×pi/w = 0.00606 s

2)Frequency = 1/T = 1/0.00606 = 165 Hz

F = 165 rev/s

3) a = w2 r = 1036.725×1036.725×0.06 = 64488 m/s2 = 6580.4g